Check List for Particle : a_{2}^{0}(1320)

March 25, 2009

Quantity  Us  PDG  EvtGen  Quantity  Us  PDG  EvtGen
PDG  115  Mass  1318.3

Code

Width 107 Gen a_20

Name

Decay  0 m.

Length

Latex Name a_{2}^{0}(1320)  Mass Gen. a_20mass Width Gen.  a_20width  Type meson
Type  normal  Con. Mass  0  Rating  ∗ ∗ ∗ ∗ ∗  Variable  fixed

Ratio

Stable  unstable  Spin  2  Charge  0  Colour  0
Lower Cut  214  Upper cut  214  PDG link  m012

• Is the description present and correct?

The a_{2}^{0}(1320) is the neutral member of the isospin triplet from the lightest tensor multiplet. The mass and width are taken from [?]. The limit on the off-shellness of the particle is set to twice the width. The decay modes are similar to those in EvtGen apart from the modelling of ωππ as ωρ, the inclusion of the η′π and γγ modes the omission of the πγ mode.

• References?

Branching Rating Outgoing Description  Decayer  EvtGen
Ratio  Particles
0.349434 ∗ ∗ ∗ ∗ ∗ ρ^{+}, π^{−}  Herwig::TensorMesonVectorPScalar::TensorVP
0.349434 ∗ ∗ ∗ ∗ ∗ ρ^{−}, π^{+}  Herwig::TensorMesonVectorPScalar::TensorVP
0.144053 ∗ ∗ ∗ ∗ ∗  π^{0}, η  Herwig::TensorMeson2PScalar::Tensor2PScalar
0.103473  ∗ ∗ ∗∗  ω, ρ^{0}  Herwig::TensorMesonVectorVector::TVV
0.024184 ∗ ∗ ∗ ∗ ∗ K^{+}, K^{−}  Herwig::TensorMeson2PScalar::Tensor2PScalar
0.024184 ∗ ∗ ∗ ∗ ∗ K^{0}, K̄^{0}  Herwig::TensorMeson2PScalar::Tensor2PScalar
0.005229  ∗ ∗ ∗∗  η′, π^{0}  Herwig::TensorMeson2PScalar::Tensor2PScalar
0.000009  ∗ ∗ ∗∗  γ, γ  Herwig::TensorMesonVectorVector::TVV

Table 1: The decay modes of the a_{2}^{0}(1320).

Total branching ratio is 1.